How can I show my 5th grader how to solve this problem in a simple way? 3 x 66= (j x 60) + (3 x 6)
Answer by ReadingBoosters(3246)

Starting with
3x66 = (jx60)+(3x6)
...
Start with the parentheses
jx60 is the same as writing 60j
3x6=18
...
So far,
3x66 = 60j + 18
..
Solve 3x66 = 198
Now,
198 = 60j + 18
..
Isolate the variable, j. Keep it balanced, by doing the same action to both sides.
198 - 18 = 60j + 18 - 18
Subtract
180 = 60j
Divide by 60 to further isolate the variable, j. Remember to keep it balanced.
180/60 = 60j/60
180/60 = j
180/60 = 3 = j
...
j=3
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com